A Treasury bond that you own at the beginning of the year is worth $1,080. During the year, it pays $44 in interest payments and ends the year valued at $1,090. What was your dollar return and percent return? (Round "Percent return" to 2 decimal places.)
Answer #1

Dollar return=(End value-Beginning value+Interest payments)

=(1090-1080+44)=$54

% return=Dollar return/Beginning value

=(54/1080)=5%